# TrIM, Triangular Input Movement Systolic Array for Convolutional Neural Networks:

### v2

@@ Abstract
In order to follow the ever-growing computational complexity and data intensity of state-of-the-art AI models, new computing paradigms are being proposed. These paradigms aim at achieving high energy efficiency, by mitigating the Von Neumann bottleneck that relates to the energy cost of moving data between the processing cores and the memory. Convolutional Neural Networks (CNNs) are susceptible to this bottleneck, given the massive data they have to manage. Systolic Arrays (SAs) are promising architectures to mitigate the data transmission cost, thanks to high data utilization of Processing Elements (PEs). These PEs continuously exchange and process data locally based on specific dataflows (like weight stationary and row stationary), in turn reducing the number of memory accesses to the main memory. In SAs, convolutions are managed either as matrix multiplications or exploiting the raster-order scan of sliding windows. However, data redundancy is a primary concern affecting area, power and energy. In this paper, we propose TrIM: a novel dataflow for SAs based on a Triangular Input Movement and compatible with CNN computing. TrIM maximizes the local input utilization, minimizes the weight data movement and solves the data redundancy problem. Furthermore, TrIM does not incur the significant on-chip memory penalty introduced by the row stationary dataflow. When compared to state-of-the-art SA dataflows the high data utilization offered by TrIM guarantees 10$\sim$ 10$\times$ 10 × less memory access. Furthermore, considering that PEs continuously overlap multiplications and accumulations, TrIM achieves high throughput (up to 81.8% higher than row stationary), other than requiring a limited number of registers (up to 15.6×15.6$\times$15.6 × fewer registers than row stationary).
@@
publicationid: pubid: This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
@@ I Introduction
Convolutional Neural Network (CNN)[2] is an example of data-intensive DNN, since it executes convolutions on multi-dimensional arrays, named feature maps, to carry out tasks like image classification[3], image segmentation[4, 5], image generation[6, 7], object detection[8, 9] and speech recognition[10, 11]. Conventionally, Central Processing Units (CPUs) and Graphics Processing Units (GPUs) manage CNNs’ workloads. However, these architectures suffer from the Von Neumann bottleneck[12], owing to the physical separation between the computing core and the memory. This significantly degrades the energy efficiency, given that data should be first fetched from an external Dynamic Random Access Memory (DRAM), then buffered on-chip, and finally processed by the computing core. For example, the normalized DRAM energy cost from a commercial 65nm process is 200$\times$ higher than the cost associated to a Multiply-Accumulation (MAC)[13], which is the elementary operation performed by the computing core. High data utilization[14] is a way to mitigate such cost, by allowing inputs, weights, or partial sums (psums) to be held on-chip as long as they need to be consumed.
@@
Systolic Arrays (SAs) are representative architectures that maximize data utilization by using an array of Processing Elements (PEs) interconnected with each other[15]. Data moves rhythmically, thus evoking the blood flow into the cardiovascular system, hence the term systolic. Conceived late 1970s[16], they have been mainly used for matrix multiplications. In recent years, the research community has put effort to allow SAs to meet the CNN’s workflow[17, 18, 19]. For instance, the Convolution to General Matrix Multiplication conversion (Conv-to-GeMM)[20] introduces data redundancy to make inputs compliant with the SAs’ dataflows. However, this reflects in higher memory requirements and, in turn, a higher number of memory accesses, thus negatively impacting area and energy. Weight Stationary (WS) based SAs are examples of architectures using Conv-to-GeMM. Inputs are moved and reused in one direction along the array, while weights are kept stationary. However, First-In-First-Out (FIFO) buffers must assist the data transfer from/to the memory, thus negatively affecting area, power and energy. The Google Tensor Processing Unit (TPU) is a WS-based SA consisting of $256\times 256$ PEs, which outperforms CPUs and GPUs by 30×30$\times$30 × in terms of energy efficiency[21]. In the Row Stationary (RS) dataflow[13], rows of inputs and weights are reused at the PE level through dedicated memory blocks, without requiring Conv-to-GeMM conversion. In this case, data redundancy is moved at the array level, where inputs are shared diagonally over multiple PEs, while weights are shared horizontally. In addition, inputs and weights circulate cycle-by-cycle inside each PE, thus reducing the energy efficiency, other than making the micro-architecture of PEs more complex. Finally, the area covered by the SA depends on the inputs’ sizes, thus making the deployment of large-scale architectures a challenge. Eyeriss[22] is the pioneer RS-based SA and consists of 168 PEs, outperforming existing SAs from 1.4×1.4$\times$1.4 × to 2.5×2.5$\times$2.5 × in terms of energy efficiency.
@@
To address the drawbacks of previous art, we propose TrIM, a new dataflow that exploits a triangular input movement to maximize input utilization. A generic SA dealing with TrIM consists of $K\times K$ PEs, where weights are kept stationary and psums are propagated vertically and finally accumulated by an extra adder tree. The TrIM-based SA is compatible with the computation requirements of Convolutional Layers (CLs). While we took inspiration from the WS dataflow to keep weights fixed throughout the computations, we considered the convolution-oriented abstraction of the RS dataflow to overcome the Conv-to-GeMM transformation. However, TrIM sits completely apart from the above referenced dataflows because it solves the data redundancy problem: differently from the baseline WS, TrIM does not require Conv-to-GeMM conversion. Also, TrIM is not penalized by the significant energy requirements associated with the on-chip memory blocks unlike the RS dataflow. Furthermore, TrIM avoids the data redundancy at the array level presented by RS. When compared to WS, TrIM exhibits one order of magnitude less memory accesses. Moreover, TrIM ensures high performance efficiency since each PE works at the peak throughput of 2 OPs/cycle, which results in 81.8% improvement of RS. Finally, the micro-architecture of PEs is fairly simple, translating into a limited number of registers, up to 15.6×15.6$\times$15.6 × lower than RS with a kernel size of 7.
@@
  • We present an analytical model for TrIM and state-of-the-art dataflows (i.e., WS and RS), including metrics like memory accesses, latency, throughput, number of registers.
@@
  • According to the design space evaluation, we discuss the key benefits offered by TrIM.
@@
The remainder of this paper is structured as follows: Section II, after introducing CNNs, provides a background about SAs and specific dataflows to manage convolutions; Section III introduces the proposed TrIM dataflow; Section IV presents an analytical model to characterize TrIM, RS and WS; using the referred model, Section V presents an in-depth design space evaluation to spotlight the advantages of TrIM over its counterparts; finally, section VI concludes the paper. In order to ease the readability of this manuscript, Table I collects all the abbreviations.
@@ II-A Convolutional Neural Networks
A CNN consists of a sequence of CLs that extract features of interest from input data[23], by emulating the behavior of human visual cortex to retrieve patterns, such as shapes and edges, from images. The feature maps (fmaps) generated from the last CL are eventually processed by Fully-Connected Layers (FCLs)[24] for classification. As shown in Fig. 1(a) Each CL performs $N$ three-dimensional convolutions between $M$ input fmaps (ifmaps), each consisting of a $H_{I}\times W_{I}$ plane, and $N$ filters, each consisting of $M$ kernels of $K_{H}\times K_{W}$ weights. In other words, each filter scans the $M$ ifmaps through sliding windows with stride $S$. In the rest of the paper, we consider $K_{H}=K_{W}=K$ and $S=1$ as in common CNNs for classification. As a result, $N$ output fmaps (ofmaps), each $H_{O}\times W_{O}$ wide, are generated, with $H_{O}=H_{I}-K+1$ and $W_{O}=W_{I}-K+1$. Each output element is named activation and follows equation (1):
@@
$O$ identifies the generic output activation, $I$ is the generic input activation and $W$ is the generic weight belonging to a kernel of one filter; $n$ iterates over the $N$ ofmaps, $h_{o}$ iterates over the $H_{O}$ rows of each ofmaps, $w_{o}$ iterates over the $W_{O}$ elements belonging to each ofmap’s row, $m$ iterates over the $M$ ifmaps, $k_{h}$ and $k_{w}$ iterate over the kernels’ rows and columns, respectively. $N$ biases can be eventually added to each output activation if required by the layer. Fig. 1(b) shows how a convolution between a $5\times 5$ ifmap and a $3\times 3$ kernel works.
@@
Figure 1: Example of Convolutional Layer (CL): (a) a set of $N$ three-dimensional filters are subjected to as many three-dimensional convolutions with $M$ ifmaps. As a result, $N$ ofmaps are generated. The portion of ifmaps highlighted in red is currently processed by the filters 0, in order to generate the red activation into the first ofmap; (b) example of two-dimensional convolution between a $5\times 5$ ifmap and a $3\times 3$ kernel. Each yellow box reports which multiplications are needed to generate the output $O_{i,j}$, with $30\leq i,j<$ 3.
@@ II-B Systolic Arrays
A SA is spatial architecture consisting of a 2-D array of PEs interconnected with each other in proper directions. Each PE usually performs a MAC operation using inputs supplied by either the main memory or neighbor PEs. With specific reference to DNNs, PEs are fed by input activations (inputs), weights and psums. High data utilization is met at two hierarchical levels: (a) at the PE level, one element among the input, weight and psum is retained in a register as long as it is required; (b) at the SA level, the other elements move rhythmically between adjacent PEs. The reused data at the PE level equips the SA with a specific stationary dataflow:
  • Weight Stationary (WS): weights are retained inside the PEs and do not move. Inputs and psum moves between adjacent PEs throughout the process.
@@ II-B1 Weight Stationary SAs
Figure 2: WS-based SA: (a) a general architecture to manage filters with $3\times 3$ kernels is reported. The different PEs’ columns manage the kernels of $N$ different filters, which in turn receive the same ifmap. Inputs move from left to right, while psums move from top to bottom. FIFOs are needed to correctly align inputs and psums over the time. The generic PE performs a MAC operation between the current input, weight and psum as well. Each data is registered to enable proper synchronization; (b) example of General Matrix-Multiplication (GeMM) between a $5\times 5$ ifmap and a $3\times 3$ kernel: the ifmap is reshaped as a $9\times 9$ matrix, where redundant data are highlighted in red, while the kernel is reshaped as a vertical $9\times 1$ array.
In a WS-based SAs, weights are preliminary fetched from the main memory and stored inside the PEs. Fig. 2(a) depicts an example of WS-based SA, where inputs are loaded from the left side and moved horizontally across the array. Differently, psums are accumulated following vertical interconnections. In order to correctly align data over time, First-In-First-Out buffers (FIFOs) for inputs and psums are placed at the left and bottom boundaries of the array, respectively. To make WS-based SAs compliant with CNNs, inputs and weights are subjected to Conv-to-GeMM[20]. The $M$ ifmaps, consisting of $H_{I}\times W_{I}$ elements each, are reshaped as a matrix having $H_{O}\times W_{O}$ rows and $M\times K\times K$ columns. Each row contains one of the three-dimensional sliding windows covered by the filters that scan the ifmaps. Filters are reshaped as $N$ column arrays, each consisting of $M\times K\times K$ entries (i.e., the number of weights of each filter). As a result, a GeMM-oriented SA consists of $M\times K\times K$ rows of $N$ PEs each. In addition, $M\times K\times K-1$ FIFOs must be adopted to ensure proper data-alignment over time. Fig 2(b) shows an example of Conv-to-GeMM between a $5\times 5$ ifmap and a $3\times 3$ kernel. Despite the above PEs’ deployment allows WS-based SAs to meet CNNs’ workloads, some drawbacks are evident: (i) Conv-to-GeMM introduces data redundancy since inputs are reshaped to guarantee the overlapping sliding windows of the original workflow. As a result, the main memory demands higher capacity, as well as higher number of memory accesses to feed the SA. These aspects negatively impact area and energy efficiency; (ii) furthermore, the larger the kernel size $K$ and/or the number of kernels $M$ and/or the number of filters $N$, the bigger the FIFOs. As a consequence, higher latency and switching activity impact the energy consumption, other than the area.
@@ II-B2 Row Stationary SAs
Figure 3: Example of RS-based SA considering a $5\times 5$ ifmap and a $3\times 3$ kernel: (a) the architecture consists of $3\times 3$ PEs, where weights’ rows are broadcast horizontally, inputs’ rows are broadcast diagonally (in the figure, different numbers and colors refer to different rows). Finally, psums are accumulated vertically. The generic PE, other than accommodating a MAC unit, hosts two memory blocks to circulate rows of inputs and weights, respectively, and a register to save the current psum. In addition, a multiplexer allows the adder to consider either the psum coming from a different PE or the internal psum; (b) an example, where the first three columns of each input row are processed. In each yellow box, which represent a specific PE, the products over time are reported. For instance, the $PE_{0,0}$ executes $1\times A$ at cycle 1, $2\times B$ at cycle 2, and $3\times C$ at cycle 3. For each cycle, the previous product is also accumulated to the current one. At the SA level, psums are accumulated vertically at cycle 4 (first row and second row of the array) and at cycle 5 (second row and third row).
@@
In a RS-based SA, the PEs are arranged as an array of $K$ rows and $H_{O}$ columns, where: (a) weights are fetched from the main memory and broadcast horizontally to all the PEs; (b) inputs are fetched and broadcast diagonally; (c) psum are accumulated vertically, by exploiting dedicated interconnections between PEs. High data utilization is handled at the PE level, despite requiring memory blocks to circulate inputs and weights. These data are typically loaded as rows of $K$ elements, in order to allow the generic PE to perform 1-D convolutions. In other words, each PE is capable to perform $K$ MACs, before forwarding the temporary psum vertically to another PE, which in the meantime has completed another 1-D convolution. The vertical connections finalize the 2-D convolution. Fig. 3(a) shows an example of RS-based SA using $3\times 3$ PEs, where a $5\times 5$ ifmap is subjected to a convolution with a $3\times 3$ kernel. The generic PE consists of a MAC unit, two memory blocks for inputs and weights, as well as control logic. Fig. 3(b) illustrates the 2-D convolutions of a $5\times 5$ ifmap, with details on the data shared among the PEs. Differently from Conv-to-GeMM, this dataflow moves data redundancy at the array level, where inputs are shared diagonally, while weights are shared horizontally. In addition, other drawbacks arise: (i) the hardware complexity of each PE is higher, considering that memory blocks (and proper control logic as well) are needed. Other than the area, this negatively impacts the energy dissipation due to the high switching activity of memory blocks to guarantee data circulation; (ii) despite inputs/weights broadcasting at the SA level enables parallel convolutions, thus the possibility to reduce the latency, this directly impacts the routability of the PEs, making the physical implementation phase challenging; (iii) finally, the area covered by the SA depends on ifmap’s size, thus limiting scalability and PEs’ utilization.

### v3

@@ Abstract
In order to follow the ever-growing computational complexity and data intensity of state-of-the-art AI models, new computing paradigms are being proposed. These paradigms aim at achieving high energy efficiency by mitigating the Von Neumann bottleneck that relates to the energy cost of moving data between the processing cores and the memory. Convolutional Neural Networks (CNNs) are susceptible to this bottleneck, given the massive data they have to manage. Systolic arrays (SAs) are promising architectures to mitigate data transmission cost, thanks to high data utilization of Processing Elements (PEs). These PEs continuously exchange and process data locally based on specific dataflows (such as weight stationary and row stationary), in turn reducing the number of memory accesses to the main memory. In SAs, convolutions are managed either as matrix multiplications or exploiting the raster-order scan of sliding windows. However, data redundancy is a primary concern affecting area, power, and energy. In this paper, we propose TrIM: a novel dataflow for SAs based on a Triangular Input Movement and compatible with CNN computing. TrIM maximizes the local input utilization, minimizes the weight data movement, and solves the data redundancy problem. Furthermore, TrIM does not incur the significant on-chip memory penalty introduced by the row stationary dataflow. When compared to state-of-the-art SA dataflows, the high data utilization offered by TrIM guarantees 10$\sim$ 10$\times$ 10 × less memory access. Furthermore, considering that PEs continuously overlap multiplications and accumulations, TrIM achieves high throughput (up to 81.8% higher than row stationary), other than requiring a limited number of registers (up to 15.6×15.6$\times$15.6 × fewer registers than row stationary).
@@ I Introduction
Convolutional Neural Network (CNN)[2] is an example of data-intensive DNN, since it executes convolutions on multi-dimensional arrays, named feature maps, to carry out tasks like image classification[3], image segmentation[4, 5], image generation[6, 7], object detection[8, 9] and speech recognition[10, 11]. In those applications, CNNs extract local features to determine the intricate relationships of input data. This capability may be further exploited in recent Generative AI models, for example in Convolutional Vision Transformers (CvTs)[12] and Stable Diffusion models[13]. Conventionally, Central Processing Units (CPUs) and Graphics Processing Units (GPUs) manage CNNs’ workloads. However, these architectures suffer from the Von Neumann bottleneck[14], owing to the physical separation between the computing core and the memory. This significantly degrades the energy efficiency, given that data should be first fetched from an external Dynamic Random Access Memory (DRAM), then buffered on-chip, and finally processed by the computing core. For example, the normalized DRAM energy cost from a commercial 65nm process is 200$\times$ higher than the cost associated to a Multiply-Accumulation (MAC)[15], which is the elementary operation performed by the computing core. High data utilization[16] is a way to mitigate such cost, by allowing inputs, weights, or partial sums (psums) to be held on-chip as long as they need to be consumed.
@@
Systolic Arrays (SAs) are representative architectures that maximize data utilization by using an array of Processing Elements (PEs) interconnected with each other[17]. Data moves rhythmically, thus evoking the blood flow into the cardiovascular system, hence the term systolic. Conceived late 1970s[18], they have been mainly used for matrix multiplications. In recent years, the research community has put effort to allow SAs to meet the CNN’s workflow[19, 20, 21]. For instance, the Convolution to General Matrix Multiplication conversion (Conv-to-GeMM)[22] introduces data redundancy to make inputs compatible with SAs’ dataflows. However, this reflects in higher memory requirements and, in turn, a higher number of memory accesses, thus negatively impacting area and energy. Weight Stationary (WS) based SAs are examples of architectures using Conv-to-GeMM. Inputs are moved and reused in one direction along the array, while weights are kept stationary. However, First-In-First-Out (FIFO) buffers must assist the data transfer from/to the memory, thus negatively affecting area, power, and energy. The Google Tensor Processing Unit (TPU) is a WS-based SA consisting of $\times 256$ 256 PEs, which outperforms CPUs and GPUs by 30×30$\times$30 × in terms of energy efficiency[23]. In the Row Stationary (RS) dataflow[15], rows of inputs and weights are reused at the PE level through dedicated memory blocks, without requiring Conv-to-GeMM conversion. In this case, data redundancy is moved at the array level, where inputs are shared diagonally over multiple PEs, while weights are shared horizontally. In addition, inputs and weights circulate cycle-by-cycle inside each PE, thus reducing the energy efficiency, besides making the micro-architecture of PEs more complex. Finally, the area covered by the SA depends on the inputs’ sizes, thus making the deployment of large-scale architectures a challenge. Eyeriss[24] is the pioneer RS-based SA and consists of 168 PEs, outperforming existing SAs from 1.4×1.4$\times$1.4 × to 2.5×2.5$\times$2.5 × in terms of energy efficiency.
@@
To address the drawbacks of previous art, we propose TrIM, a new dataflow that exploits a triangular input movement to maximize input utilization. A generic SA dealing with TrIM consists of $K\times K$ PEs, where weights are kept stationary and psums are propagated vertically and finally accumulated by an extra adder tree. The TrIM-based SA is compatible with the computational requirements of Convolutional Layers (CLs). While we took inspiration from the WS dataflow to keep weights fixed throughout the computations, we considered the convolution-oriented abstraction of the RS dataflow to overcome the Conv-to-GeMM transformation. However, TrIM sits completely apart from the above referenced dataflows because it solves the data redundancy problem: differently from the baseline WS, TrIM does not require Conv-to-GeMM conversion. Also, TrIM is not penalized by the significant energy requirements associated with the on-chip memory blocks unlike the RS dataflow. Furthermore, TrIM avoids the data redundancy at the array level presented by RS. When compared to WS, TrIM exhibits one order of magnitude less memory accesses. Moreover, TrIM ensures high performance efficiency, as each PE works at the peak throughput of 2 OPs/cycle, which results in 81.8% improvement of RS. Finally, the micro-architecture of PEs is fairly simple, translating into a limited number of registers, up to 15.6×15.6$\times$15.6 × lower than RS with a kernel size of 7.
@@
  • We present an analytical model for TrIM and state-of-the-art dataflows (i.e., WS and RS), including metrics like memory accesses, latency, throughput, and number of registers.
@@
  • Based on the design space evaluation, we discuss the key benefits offered by TrIM. We also provide a perspective on the exploitation of TrIM in system-level implementations for state-of-the-art AI models.
@@
The remainder of this paper is structured as follows: Section II, after introducing CNNs, provides a background about SAs and specific dataflows to manage convolutions; Section III introduces the proposed TrIM dataflow; Section IV presents an analytical model to characterize TrIM, RS and WS; using the referred model, Section V presents an in-depth design space evaluation to spotlight the advantages of TrIM over its counterparts; Section VI provides a perspective on how the TrIM concept can be applied in recent AI models, how can be hardware implemented, and how it compares to the most recent dataflow optimizations; finally, section VII concludes the paper. In order to facilitate the readability of this manuscript, Table I collects all the abbreviations.
@@ II-A Convolutional Neural Networks
A CNN consists of a sequence of CLs that extract features of interest from input data[25], by emulating the behavior of the human visual cortex to retrieve patterns, such as shapes and edges, from images. The feature maps (fmaps) generated from the last CL are eventually processed by Fully-Connected Layers (FCLs)[26] for classification. As shown in Fig. 1(a) Each CL performs $N$ three-dimensional convolutions between $M$ input fmaps (ifmaps), each consisting of a $H_{I}\times W_{I}$ plane, and $N$ three-dimensional filters. Each filter volume is made of $M$ matrices that are defined as kernels, whose weights are responsible for specific features extraction. Each filter scans the $M$ ifmaps through sliding windows with stride $S$. In the rest of the paper, we consider $K_{H}=K_{W}=K$ and $S=1$ as in common CNNs for classification. As a result, $N$ output fmaps (ofmaps), each $H_{O}\times W_{O}$ wide, are generated, with $H_{O}=H_{I}-K+1$ and $W_{O}=W_{I}-K+1$. Each output element is named activation and follows equation (1):
@@
$O$ identifies the generic output activation, $I$ is the generic input activation and $W$ is the generic weight belonging to a kernel of one filter; $n$ iterates over the $N$ ofmaps, $h_{o}$ iterates over the $H_{O}$ rows of each ofmaps, $w_{o}$ iterates over the $W_{O}$ elements belonging to each ofmap’s row, $m$ iterates over the $M$ ifmaps, $k_{h}$ and $k_{w}$ iterate over the kernels’ rows and columns, respectively. $N$ biases can be eventually added to each output activation if required by the layer. Fig. 1(b) shows how a convolution between a $\times 5$ 5 ifmap and a $\times 3$ 3 kernel works.
@@
Figure 1: Example of Convolutional Layer (CL): (a) a set of $N$ three-dimensional filters are subjected to as many three-dimensional convolutions with $M$ ifmaps. As a result, $N$ ofmaps are generated. The portion of ifmaps highlighted in red is currently processed by the filters 0, in order to generate the red activation into the first ofmap; (b) example of two-dimensional convolution between a $\times 5$ 5 ifmap and a $\times 3$ 3 kernel. Each yellow box reports which multiplications are needed to generate the output $O_{i,j}$, with $0\leq i,j<3$.
@@ II-B Systolic Arrays
A SA is spatial architecture consisting of a 2-D array of PEs interconnected with each other in proper directions. Each PE usually performs a MAC operation using inputs supplied by either the main memory or neighbor PEs. With specific reference to DNNs, PEs are fed by input activations (inputs), weights and psums. High data utilization is met at two hierarchical levels: (a) at the PE level, one element among the input, weight and psum is retained in a register as long as it is required; (b) at the SA level, the other elements move rhythmically between adjacent PEs. The reused data at the PE level equip the SA with a specific stationary dataflow:
  • Weight Stationary (WS): weights are retained inside the PEs and do not move. Inputs and psum move between adjacent PEs throughout the process.
@@ II-B1 Weight Stationary SAs
Figure 2: WS-based SA: (a) a general architecture to manage filters with $\times 3$ 3 kernels is reported. The different PEs’ columns manage the kernels of $N$ different filters, which in turn receive the same ifmap. Inputs move from left to right, while psums move from top to bottom. FIFOs are needed to correctly align inputs and psums over the time. The generic PE performs a MAC operation between the current input, weight and psum as well. Each data is registered to enable proper synchronization; (b) example of General Matrix-Multiplication (GeMM) between a $\times 5$ 5 ifmap and a $\times 3$ 3 kernel: the ifmap is reshaped as a $\times 9$ 9 matrix, where redundant data are highlighted in red, while the kernel is reshaped as a vertical 9×19$\times$ 19 × 1 array.
In a WS-based SAs, weights are preliminary fetched from the main memory and stored inside the PEs. Fig. 2(a) depicts an example of WS-based SA, where inputs are loaded from the left side and moved horizontally across the array. Differently, psums are accumulated following vertical interconnections. In order to correctly align data over time, First-In-First-Out buffers (FIFOs) for inputs and psums are placed at the left and bottom boundaries of the array, respectively. To make WS-based SAs compatible with CNNs, inputs and weights are subjected to Conv-to-GeMM[22]. The $M$ ifmaps, consisting of $H_{I}\times W_{I}$ elements each, are reshaped as a matrix having $H_{O}\times W_{O}$ rows and $M\times K\times K$ columns. Each row contains one of the three-dimensional sliding windows covered by the filters that scan the ifmaps. Filters are reshaped as $N$ column arrays, each consisting of $M\times K\times K$ entries (i.e., the number of weights of each filter). As a result, a GeMM-oriented SA consists of $M\times K\times K$ rows of $N$ PEs each. In addition, $M\times K\times K-1$ FIFOs must be adopted to ensure proper data-alignment over time. Fig. 2(b) shows an example of Conv-to-GeMM between a $\times 5$ 5 ifmap and a $\times 3$ 3 kernel. Despite the above PEs’ deployment allows WS-based SAs to meet CNNs’ workloads, some drawbacks are evident: (i) Conv-to-GeMM introduces data redundancy since inputs are reshaped to guarantee the overlapping sliding windows of the original workflow. As a result, the main memory demands higher capacity, as well as higher number of memory accesses to feed the SA. These aspects negatively impact area and energy efficiency; (ii) furthermore, the larger the kernel size $K$ and/or the number of kernels $M$ and/or the number of filters $N$, the bigger the FIFOs. As a consequence, higher latency and switching activity impact the energy consumption, other than the area.
@@ II-B2 Row Stationary SAs
Figure 3: Example of RS-based SA considering a $\times 5$ 5 ifmap and a $\times 3$ 3 kernel: (a) the architecture consists of $\times 3$ 3 PEs, where weights’ rows are broadcast horizontally, inputs’ rows are broadcast diagonally (in the figure, different numbers and colors refer to different rows). Finally, psums are accumulated vertically. The generic PE, other than accommodating a MAC unit, hosts two memory blocks to circulate rows of inputs and weights, respectively, and a register to save the current psum. In addition, a multiplexer allows the adder to consider either the psum coming from a different PE or the internal psum; (b) an example, where the first three columns of each input row are processed. In each yellow box, which represent a specific PE, the products over time are reported. For instance, the $PE_{0,0}$ executes $1\times A$ at cycle 1, $2\times B$ at cycle 2, and $3\times C$ at cycle 3. For each cycle, the previous product is also accumulated to the current one. At the SA level, psums are accumulated vertically at cycle 4 (first row and second row of the array) and at cycle 5 (second row and third row).
@@
In a RS-based SA, the PEs are arranged as an array of $K$ rows and $H_{O}$ columns, where: (a) weights are fetched from the main memory and broadcast horizontally to all the PEs; (b) inputs are fetched and broadcast diagonally; (c) psum are accumulated vertically, by exploiting dedicated interconnections between PEs. High data utilization is handled at the PE level, despite requiring memory blocks to circulate inputs and weights. These data are typically loaded as rows of $K$ elements, in order to allow the generic PE to perform 1-D convolutions. In other words, each PE is capable to perform $K$ MACs, before forwarding the temporary psum vertically to another PE, which in the meantime has completed another 1-D convolution. The vertical connections finalize the 2-D convolution. Fig. 3(a) shows an example of RS-based SA using $\times 3$ 3 PEs, where a $\times 5$ 5 ifmap is subjected to a convolution with a $\times 3$ 3 kernel. The generic PE consists of a MAC unit, two memory blocks for inputs and weights, as well as control logic. Fig. 3(b) illustrates the 2-D convolutions of a $\times 5$ 5 ifmap, with details on the data shared among the PEs. Differently from Conv-to-GeMM, this dataflow moves data redundancy at the array level, where inputs are shared diagonally, while weights are shared horizontally. In addition, other drawbacks arise: (i) the hardware complexity of each PE is higher, considering that memory blocks (and proper control logic as well) are needed. Other than the area, this negatively impacts the energy dissipation due to the high switching activity of memory blocks to guarantee data circulation; (ii) despite inputs/weights broadcasting at the SA level enables parallel convolutions, thus the possibility to reduce the latency, this directly impacts the routability of the PEs, making the physical implementation phase challenging; (iii) finally, the area covered by the SA depends on ifmap’s size, thus limiting scalability and PEs’ utilization.